OpenStudy (anonymous):

graph -3x^2+12y^2=849 what are the domain and range?

OpenStudy (anonymous):

\[\frac{ 12y^2 }{ 849 }-\frac{ 3x^2 }{ 849 }=1\] \[\frac{ y^2 }{ \frac{ 849 }{12 } }-\frac{ x^2 }{\frac{ 849 }{3 } }=1\] \[\frac{ y^2 }{\frac{ 283 }{4 } }-\frac{ x^2 }{ 283 }=1\] it is of the type \[\frac{ y ^{2} }{a^2 }-\frac{ x^2 }{ b^2 }=1\] which is a hyperbola with axis on y-axis.

OpenStudy (anonymous):

those are my choices.